### v1

# Deformable Beta Splatting

Abstract.
3D Gaussian Splatting (3DGS) has advanced radiance field reconstruction by enabling real-time rendering. However, its reliance on Gaussian kernels for geometry and low-order Spherical Harmonics (SH) for color encoding limits its ability to capture complex geometries and diverse colors. We introduce Deformable Beta Splatting (DBS), a deformable and compact approach that enhances both geometry and color representation. DBS replaces Gaussian kernels with deformable Beta Kernels, which offer bounded support and adaptive frequency control to capture fine geometric details with higher fidelity while achieving better memory efficiency. In addition, we extended the Beta Kernel to color encoding, which facilitates improved representation of diffuse and specular components, yielding superior results compared to SH-based methods. Furthermore, Unlike prior densification techniques that depend on Gaussian properties, we mathematically prove that adjusting regularized opacity alone ensures distribution-preserved Markov chain Monte Carlo (MCMC), independent of the splatting kernel type. Experimental results demonstrate that DBS achieves state-of-the-art visual quality while utilizing only 45% of the parameters and rendering 1.5x faster than 3DGS-based methods. Notably, for the first time, splatting-based methods outperform state-of-the-art Neural Radiance Fields, highlighting the superior performance and efficiency of DBS for real-time radiance field rendering.
journal: TOG
Figure 1. Superior scene reconstruction by Beta Splatting. Our deformable Beta Kernels achieve high-fidelity geometry and appearance representation (a), while naturally enabling the decoupling of surface geometry from texture details (b) and the separation of specular and diffuse lighting components (c).
1. Introduction
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021; Barron et al., 2021; Müller et al., 2022; Barron et al., 2023) have demonstrated exceptional capabilities in 3D scene representation and novel view synthesis but are hindered by high computational costs and limited applicability in real-time scenarios due to their reliance on neural network-based encoding and dense volumetric ray-marching. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduced an explicit representation paradigm that shifts from volumetric ray-marching to point-based splatting. By representing scenes as a set of 3D Gaussian primitives, 3DGS significantly reduces computational overhead, achieving real-time radiance field rendering while maintaining visual quality comparable to earlier NeRF methods. This explicit representation eliminates the need for heavy neural networks and enables direct manipulation of scene geometry and real-time rendering, making 3DGS ideal for applications such as immersive VR exploration (Jiang et al., 2024; Franke et al., 2024; Yu et al., 2024), dynamic scene representation (Yang et al., 2023; Luiten et al., 2023; Wu et al., 2024; Yang et al., 2024b; Guo et al., 2024), 3D content creation (Tang et al., 2024; Yi et al., 2024a; Chung et al., 2023; Yi et al., 2024b), mesh extraction (Guédon and Lepetit, 2023; Huang et al., 2024; Liu et al., 2024; Dai et al., 2024), semantic field distillation (Shi et al., 2023; Zhou et al., 2024; Qin et al., 2024; Qiu et al., 2024; Ye et al., 2024a), and cinematic production (Wang et al., 2024). However, while 3DGS excels in efficiency, its rendering quality still falls short of the latest NeRF advancements (Barron et al., 2023), particularly in capturing fine details and modeling view-dependent colors. To address this gap, a range of extensions has emerged, focusing on improving its visual quality through anti-alising (Yu et al., 2023; Yan et al., 2024; Liang et al., 2024; Song et al., 2024), deblurring (Lee et al., 2024; Peng et al., 2024; Zhao et al., 2024; Seiskari et al., 2024; Oh et al., 2024), appearance modeling (Xu et al., 2024; Svitov et al., 2024; Weiss and Bradley, 2024; Yang et al., 2024a), and revisting densification (Lu et al., 2023; Niemeyer et al., 2024; Zhang et al., 2024; Liu et al., 2024; Bulò et al., 2024; Kheradmand et al., 2024). Despite these efforts, 3DGS-based methods have yet to surpass the state-of-the-art NeRF methods (Barron et al., 2023).
The limitations of 3DGS primarily arise from the constraints of its foundational components. The Gaussian kernel, while enabling real-time rendering through its explicit representation, inherently produces smooth and blurred results. This smoothing effect hinders its ability to accurately depict flat surface geometries and sharp edges, resulting in a loss of fine detail and diminished realism in rendered scenes. Additionally, the Gaussian kernel’s long-tailed and unbounded nature necessitates a hard-coded cut-off for efficient rendering. This workaround introduces cut-off artifacts and further compromises the accurate representation of fine geometric details. Furthermore, 3DGS employs low–order Spherical Harmonics (SH) to encode view-dependent color information. Although SH can model lighting variations, its parameter count increases polynomially with the SH level, limiting practical use in real-time applications. As a result, 3DGS relies on lower SH levels, which inherently produce smooth and overly simplified color transitions. This prevents the accurate modeling of specular highlights and fails to decouple diffuse and specular components, resulting in suboptimal representations of complex lighting interactions. Notably, both Gaussian kernel and SH are fixed-function mappings, restricting the adaptability required to handle varying scene complexities and lighting conditions, further compromising visual fidelity.
To address these shortcomings, we propose Deformable Beta Splatting (DBS), a novel framework designed to enhance both geometric and color representation in radiance field rendering. At the core of DBS is the Beta Kernel, inspired by the Beta distribution, which provides bounded support and flexible shape control. This kernel adapts naturally to diverse scene complexities, enabling accurate representation of flat surfaces, sharp edges, and smooth regions with high fidelity. For color representation, we introduce the Spherical Beta (SB) function for view-dependent color encoding. By decoupling diffuse and specular components, SB effectively models sharp specular reflections and captures high-frequency lighting effects. Compared to spherical harmonics (SH) of degree 3, SB achieves superior performance using only 31% of the parameters, significantly improving model efficiency. Finally, we revisit and adapt Markov Chain Monte Carlo (MCMC) principles from 3DGS-MCMC (Kheradmand et al., 2024) into the Deformable Beta Splatting framework to address unique challenges posed by deformable kernels. Specifically, the adaptive density control in 3DGS and the probability-preserved densification strategies in 3DGS-MCMC were designed for Gaussian kernels, making their direct application to deformable Beta kernels nontrivial. To overcome this, we mathematically prove that by regularizing opacity, adjusting opacity alone is valid for distribution-preserved densification, regardless of the number of densifications or the choice of splatting kernel. This results in a kernel-agnostic MCMC strategy, eliminating the need for complex heuristics and scale recomputation, thereby streamlining both the optimization and densification processes. Additionally, we redefine the noise function used in MCMC with the Beta kernel to maintain mathematical consistency and compatibility with opacity constraints.
The comprehensive experimental results demonstrate that our approach consistently outperforms both 3DGS-based and NeRF-based methods in visual quality across all benchmarks. Furthermore, our method utilizes only 45% of the parameters required by 3DGS and renders 1.5x faster than Gaussian Splatting-based methods. It is worthy noting that, for the first time, splatting-based methods surpass state-of-the-art NeRF, demonstrating DBS’s superior performance and efficiency in real-time radiance field rendering.
To summarize, the main contributions of our paper are as follows:
  1. (1)
    Beta Kernel for Geometry Representation: A bounded, flexible kernel that improves geometric fidelity by accurately capturing flat surfaces, sharp edges, and smooth regions.
  2. (2)
    Spherical Beta for Color Encoding: An efficient view-dependent color model that decouples diffuse and specular components, enabling sharp specular highlights with fewer parameters.
  3. (3)
    Kernel-Agnostic MCMC for optimization: A distribution-preserved densification leveraging regularized opacity and a redefined Beta noise function, together enhancing optimization stability and efficiency independent of kernel shape.
2. Related Work
This section reviews key approaches in novel view synthesis, including implicit NeRFs and explicit 3DGS, highlighting how our work builds on these advancements and contributes to high-fidelity, real-time 3D rendering.
2.1. Neural Radiance Fields
The introduction of Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) revolutionized 3D scene representation and novel view synthesis. By encoding a continuous volumetric scene within a multilayer perceptron (MLP), NeRFs jointly learn scene density and view-dependent color, enabling photorealistic novel view synthesis through differentiable volumetric rendering.
Since their inception, NeRFs have spurred extensive research to address challenges in convergence speed, computational efficiency, and scalability. Early advancements, such as FastNeRF (Garbin et al., 2021) and PlenOctrees (Yu et al., 2021), improved efficiency through caching strategies and spatial acceleration structures. Methods like NeRF++(Zhang et al., 2020) and DeRF(Rebain et al., 2021) introduced optimized scene decompositions and novel optimization techniques, further accelerating training and inference. Additionally, techniques such as TensoRF (Chen et al., 2022) and DVGO (Sun et al., 2022) leveraged tensor factorization and voxel grids to reduce memory requirements and computation time. To address scalability, methods like Mip-NeRF 360 (Barron et al., 2021), Block-NeRF (Tancik et al., 2022), and UrbanNeRF (Rematas et al., 2022) introduced scene partitioning and multiscale representations for handling large and unbounded scenes. Meanwhile, approaches such as NeRF-W (Martin-Brualla et al., 2021) disentangled appearance variations, and Ref-NeRF (Verbin et al., 2022) extended NeRFs to handle reflective and refractive effects, advancing realism and robustness. Additional advancements, such as Instant-NGP (Müller et al., 2022) and Zip-NeRF (Barron et al., 2023), accelerated NeRF training by introducing a multi-resolution hash encoding and addressed aliasing issues inherent in grid-based models by incorporating anti-aliasing techniques, continuously pushing the boundaries of efficiency and quality in neural scene representation.
High-fidelity NeRF rendering requires dense ray sampling and extensive network evaluations, hindering real-time deployment. Its implicit scene representation also complicates integration with game engines and industrial pipelines, which prefer explicit geometry. These challenges drive the exploration of compact representations, explicit models, and novel optimization techniques to enable interactive, practical applications.
Figure 2. Unlike the fixed Gaussian Kernel, Deformable Beta Kernel adapts its shape to capture fine geometric and texture details. This figure shows how the kernel varies by different $b$ values as in Equation 3. Negative $b$ generates flatter surface but sharper cutoffs for learning solid support and sharp geometry, while positive $b$ generates sharper peaks for learning high frequency details. When $b=0$, the beta kernel is almost identical to Gaussian (domain scaled by $3\sigma$)
2.2. 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) offers a compelling alternative to NeRF by representing scenes as collections of anisotropic Gaussian ellipsoids that capture local properties such as position, density, and view-dependent appearance. Combining point-based rendering(Zwicker et al., 2001) with modern machine learning strategy, it enables flexible and efficient scene representation. However, despite its computational efficiency, 3DGS lags behind NeRF in photorealism, struggling with fine details, aliasing artifacts, and complex view-dependent effects.
Research has addressed these shortcomings across several fronts. Anti-aliasing methods such as Mip-Splatting (Yu et al., 2023), multi-scale approaches (Yan et al., 2024), analytic approximations (Liang et al., 2024), and scale-adaptive filtering (Song et al., 2024) have been proposed to reduce jagged edges and shimmering, thereby smoothing the rendered output. Deblurring techniques have also been explored to enhance clarity. Methods leveraging small MLPs (Lee et al., 2024), event camera data (Peng et al., 2024), joint optimization of Gaussian parameters with camera trajectory recovery (Zhao et al., 2024; Seiskari et al., 2024; Oh et al., 2024) have improved the sharpness of reconstructions from blurred inputs. These approaches integrate motion models and depth-aware filters into the 3DGS pipeline to refine both geometry and appearance.
Beyond addressing aliasing and blurring, splatting kernel modeling has emerged as an attractive direction for improving rendering quality and efficiency. 2DGS introduces a robust 2D Gaussian splatting framework that simplifies computation while producing better geometries via screen-space rendering (Huang et al., 2024). Li et al. proposed 3D Half Gaussian Splatting (3D-HGS), which optimizes the Gaussian representation by restricting it to half-space for better modeling of occlusions and interactions in complex 3D scenes (Li et al., 2024). Chen et al. developed Linear Kernel Splatting, which replaces Gaussian kernels with linear ones, enabling faster computation and maintaining high-fidelity rendering  (Chen et al., 2024). Despite advancements in previous kernel designs, their limited deformability hinders accurate representation of flat surfaces and sharp edges. In contrast, our Beta Kernel offers flexible shape control, adapting to diverse geometries and improving representation fidelity and rendering quality.
Additionally, appearance modeling has been a crucial area for improving the fidelity of 3DGS renderings. Xu et al. introduced Texture-GS, disentangling appearance from geometry by mapping 2D textures onto 3D surfaces via an MLP for high-fidelity editing (Xu et al., 2024). Svitov et al. proposed BBSplat, using textured planar primitives with learnable RGB textures and alpha maps, achieving over 1200 FPS (Svitov et al., 2024). Weiss and Bradley developed Gaussian Billboards, unifying 2D Gaussian Splatting with UV-parameterized textures for expressive scene representations (Weiss and Bradley, 2024). Yang et al. introduced Spec-Gaussian, leveraging an MLP to model anisotropic spherical Gaussians for complex lighting (Yang et al., 2024a). While existing methods focus on either speed (via billboard textures) or fidelity (via MLPs, with slower rendering), we propose Spherical Beta, which explicitly models appearance by decoupling diffuse and specular components, inspired by the Phong Reflection Model (Phong, 1975).
Revisiting optimization and densification methods has proven crucial for improving 3DGS. Techniques such as Scaffold-GS (Lu et al., 2023), RadSplat (Niemeyer et al., 2024), Pixel-GS (Zhang et al., 2024), and AtomGS (Liu et al., 2024) focus on efficient Gaussian placement and density control. Notably, 3DGS-MCMC (Kheradmand et al., 2024) introduces stochastic processes to refine the sampling and placement of Gaussians, enhancing both rendering quality and speed. These strategies ensure effective capture of scene details while balancing quality with computational cost. Building on these advancements, we extended 3DGS-MCMC principles and designed a Kernel-agnostic densification strategy that stabilizes the optimization process while ensuring spatial distribution consistency.
3. Deformable Beta Splatting
The primary bottleneck in Gaussian Splatting–based approaches lies in their reliance on a fixed, smooth splatting kernel and low-order SH encoding. To overcome this limitation, we first derive a deformable Beta Kernel function. We then illustrate how this flexible kernel can be applied to both geometric splatting and color encoding, enabling the modeling of varying geometric complexities and lighting frequencies, which enhances the overall scene representation. Finally, we address the unique challenges introduced by employing a deformable kernel and detail how our optimization strategy effectively manages these issues.
3.1. The Beta Kernel
The Beta Kernel is derived from the Beta distribution (Johnson et al., 1995), defined as a power of the input $x$ and its reflection $(1-x)$:
(1) $f(x;\alpha,\beta)=\frac{1}{\mathbf{B(\alpha,\beta)}}x^{\alpha-1}(1-x)^{\beta-1},\quad x\in[0,1],$
where $\mathbf{B}$ is the Beta Function for normalization, and $\alpha,\beta>0$ are shape-controlling parameters.
To obtain a bell-like shape that promotes smooth 3D view consistency, we fix $\alpha=1$, focusing solely on the reflection component. By removing the normalization term, we simplify the Beta distribution to the Beta Kernel:
(2) $\mathcal{B}(x;\beta)=(1-x)^{\beta},\quad x\in[0,1],\beta\in(0,\infty).$
The Beta Kernel is a power function over the closed domain $x\in[0,1]$, inherently compact and bounded, with the parameter $\beta$ introducing adaptive shape control. It is worth noting that not every function maintains multi-view consistency or qualifies as a valid splatting kernel. We have employed the inverse Abel transform to prove that the Beta Kernel meets these criteria (please refer to Appendix A for details).
[…]
As a result, we parameterize the Beta Kernel as:
(3) $\mathcal{B}(x;b)=(1-x)^{\beta(b)},\quad\beta(b)=4e^{b},\quad x\in[0,1],\quad b\in\mathbb{R}.$
This parameterization ensures that the Beta Kernel begins with a Gaussian-like shape when $b$ is initialized to zero and can adapt its shape during optimization. It eliminates the need for manual fine-tuning of initial $\beta$ values, effectively making the Beta Kernel a more flexible superset of the Gaussian Kernel.
3.2. 3D Ellipsoidal Beta Primitive
We define a 3D Ellipsoidal Beta Primitive as a set of parameters:
(4) $B=\{\bm{\mu},o,\bm{q},\bm{s},b,\bm{f}\}$
where the geometry parameters $\bm{\mu},o,\bm{q},\text{and }\bm{s}$ mirror those in 3DGS. Specifically, $\bm{\mu}\in\mathbb{R}^{3}$ denotes the center of the primitive in 3D space, $o\in[0,1]$ represents opacity, $\bm{q}\in[-1,1]^{4}$ defines the rotation through a quaternion, and $\bm{s}\in[0,\infty)^{3}$ determines the scale of the ellipsoidal shape.
The key innovation in our approach is the introduction of the parameter $b\in\mathbb{R}$, which controls the shape of the Beta Kernel. As illustrated in Figure 2, varying $b$ deforms the kernel shape, with $b=0$ yielding a function nearly identical to a Gaussian. This additional parameter provides adaptive shape control with minimal overhead, allowing the kernel to dynamically adjust from a Gaussian-like function to forms that can capture sharper edges and flat surfaces as optimization progresses.
For appearance modeling, each primitive carries a feature vector $f\in\mathbb{R}^{d}$ that encodes view-dependent color information. This vector is crucial for representing complex lighting interactions and will be discussed in detail in the next subsection.
The rendering process for a pixel $\bm{x}\in\mathbb{R}^{2}$ unfolds as follows: each 3D Ellipsoidal Beta Primitive is projected into the 2D image plane using the viewing transformation $\bm{W}$. This results in a 2D projected center $\bm{\mu}^{\prime}\in\mathbb{R}^{2}$ and a corresponding 2D covariance matrix $\bm{\Sigma}^{\prime}=\bm{JW\Sigma W^{\top}J^{\top}}$, where $\Sigma=RSS^{\top}R^{\top}$ with $R$ and $S$ derived from the quaternion and scale, respectively. We then compute the distance between the pixel and each primitive center:
(5) $r_{i}(\bm{x})=\sqrt{(\bm{x}-\bm{\mu}_{i}^{\prime})^{\top}\bm{\Sigma}_{i}^{\prime^{-1}}(\bm{x}-\bm{\mu}_{i}^{\prime})}.$
After sorting overlapping primitives by depth relative to the camera, denoted by the ordered set $\mathcal{N}=\{B_{1},\dots,B_{N}\}$, we composite their contributions to the final pixel color $\bm{C}(\bm{x})$ using the Beta Kernel:
(6) $\bm{C}(\bm{x})=\sum_{i=1}^{N}c_{i}o_{i}\mathcal{B}(r_{i}(\bm{x})^{2};b_{i})\prod_{j=1}^{i-1}\left(1-o_{j}\mathcal{B}(r_{j}(\bm{x})^{2};b_{j})\right)$
Figure 3. Analogous to the Phong Reflection Model, learnable spherical betas can effectively capture specular highlights with varying sharpness. Smaller $b$ values result in broader reflections, while higher $b$ values correspond to sharper specular highlights.
3.3. Spherical Beta
3DGS employs Spherical Harmonics (SH) for view-dependent color encoding. For a SH of degree $N$, the feature dimension scales as $3(N+1)^{2}$, resulting in a quadratic increase in parameters. To achieve real-time performance in practical applications, 3DGS is constrained to using low-order SH (e.g.,$N=3$), which only provide smooth view-dependent colors and struggle to model sharp specular highlights effectively. To overcome these issues, we introduce Spherical Beta function, inspired by the Phong Reflection Model (Phong, 1975):
(7) $c(\hat{V})=A_{m}+\sum_{m\in\mathcal{M}}\left[D_{m}+(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}\right],$
where $A_{m}$ is the ambient term, $D_{m}$ and $(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}$ models the diffuse and specular components for light source $m$. Here, $\hat{R}_{m}$ is the normalized reflection direction, $\hat{V}$ is the normalized view direction, $\alpha_{m}$ is the shininess coefficient controlling specularity, and $c_{m}$ is the color of light source. We reorganize the equation by separating view-independent and view-dependent parts and combining the view-independent ambient and diffuse terms, yielding:
(8) $c(\hat{V})=c_{0}+\sum_{m\in\mathcal{M}}(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}.$
Next, we apply the Beta Kernel to express the specular term:
(9) $c(\hat{V})=c_{0}+\sum_{m\in\mathcal{M}}\mathcal{B}(1-\hat{R}_{m}\cdot\hat{V};b_{m})c_{m}$
where $c_{0}$ represents the diffuse or base color for the Beta primitive. For each 3D Beta primitive, we introduce $M=|\mathcal{M}|$ light sources. For each light source $m\in\mathcal{M}$, it is paramertized by reflection direction $\hat{R}_{m}$, color $c_{m}$, and shininess $\beta_{m}$. We can observe that its feature dimension scales as $3+6M$, growing linearly with the number of lights and making it more efficient than SH.
Spherical Beta shares similarities with Spherical Gaussian in modeling specularities but offers improvements. Unlike Spherical Gaussian, which have unbounded support and require truncation—leading to radiance discontinuities when view and reflection directions are orthogonal—the Beta Kernel is inherently bounded. Its input range is confined to $[0,1]$, ensuring that radiance naturally drops to zero when $\hat{R}_{m}\cdot\hat{V}=0$. This eliminates artifacts associated with cutoff truncation and maintains continuity in radiance.
Figure 3 illustrates the specular highlights and corresponding reflection blobs for different $b_{m}$ values, demonstrating the capability of Spherical Beta to model both high and low frequency specularities. By integrating bounded support and adaptive shininess, Spherical Beta enhances the ability to model complex lighting interactions and specular effects with fewer parameters, thereby improving both visual fidelity and computational efficiency.
Figure 4. Geometry Decomposition: our designed Beta Kernel provides the capability to decompose scene geometry into fundamental structures and intricate details, such as high-frequency textures and fine surface variations.
Figure 5. Qualitative Comparison: Leveraging our Deformable Beta Kernel, Spherical Beta color encoding, and Kernel-Agnostic MCMC optimization process, the Deformable Beta Splatting framework qualitatively outperforms the implicit SOTA method (Zip-NeRF) and the explicit SOTA method (3DGS-MCMC).
3.4. Kernel-Agnostic Markov Chain Monte Carlo
One of the key innovations of 3DGS is its optimization with densification strategy, which utilizes heuristic gradient thresholds to clone, split, and prune primitives. Additionally, 3DGS-MCMC redefines Gaussian Splatting optimization as a Markov Chain Monte Carlo (MCMC) process by treating each primitive’s opacity as a probability. This approach identifies ”dead” primitives—those with opacity below a pruning threshold—and respawns them by relocating to live Gaussians through multinomial sampling based on their opacity values. By adjusting both opacity and scale, this method preserves the underlying probability distribution during densification, thereby stabilizing the process. Moreover, the introduction of positional noise encourages exploration and prevents overfitting.
However, these heuristic strategies and scale adjustments are inherently tied to the properties of the Gaussian function, making them difficult to directly apply to arbitrary kernels.
To address this limitation, we present Kernel-Agnostic MCMC, building upon the framework established by 3DGS-MCMC (Kheradmand et al., 2024). We first detail the optimization process and then revisit distribution-preserved densification.
Optimization
To encourage exploration and prevent overfitting, we adopt a noise term $\epsilon$ that perturbs the positions of primitives during optimization. The primitive positions $\bm{\mu}$ are then updated as
(10) $\bm{\mu}\leftarrow\bm{\mu}-\lambda_{\text{lr}}\cdot\nabla_{\bm{\mu}}\mathcal{L}+\lambda_{\epsilon}\cdot\epsilon,\quad\epsilon=\lambda_{\text{lr}}\cdot\mathcal{B}(o_{i};b^{\prime})\cdot\Sigma_{\eta}.$
Note that we replace the logit noise function from 3DGS-MCMC with Beta Kernel function for a more compact and well-defined noise function. Here, $\lambda_{\epsilon}$ controls the impact of the noise term, ensuring that the primitive is not only optimized through backpropagation of the loss function but also explores alternative solutions for better representation. $\lambda_{\text{lr}}$ represents the learning rate, and $\Sigma_{\eta}$ is the sampled noise. $\mathcal{B}(o_{i};b^{\prime})$ is the Beta Kernel component, which scales the noise based on the opacity of each primitive, with $b^{\prime}=\ln(25)$ to resemble the fall-off behavior of the original logit function.
To train the Beta Primitives, we employ a composite loss function:
(11) $\mathcal{L}=(1-\lambda_{\text{SSIM}})\mathcal{L}_{1}+\lambda_{\text{SSIM}}\mathcal{L}_{\text{SSIM}}+\lambda_{o}\sum_{i}|o_{i}|+\lambda_{\Sigma}\sum_{i}\sum_{j}|\sqrt{\text{eig}_{j}(\Sigma_{i})}|,$
where $\mathcal{L}_{1}$ is the pixel-wise L1 loss that enforces accurate color reproduction in the rendered image, and $\mathcal{L}_{\text{SSIM}}$ is the Structural Similarity Index Measure (SSIM) loss that promotes structural consistency across the scene. The hyperparameters $\lambda_{\text{SSIM}},\lambda_{o},$ and $\lambda_{\Sigma}$ balance the contributions of each loss term.
The term $\lambda_{o}\sum_{i}|o_{i}|$ acts as an opacity regularizer to ensure the kernel-agnostic densification and encourage primitive to explore. The eigenvalues $\text{eig}_{j}(\Sigma_{i})$ of the covariance matrix $\Sigma_{i}$ for the $i$-th primitive are used in the regularization term $\lambda_{\Sigma}\sum_{i}\sum_{j}|\sqrt{\text{eig}_{j}(\Sigma_{i})}|$, which together with opacity regularizer encourage primitives to disappear and respawn.
Densification
Here, we mathematically demonstrate that regularizing opacity and adjusting it alone are both effective and valid, regardless of the number of clones or the chosen splatting kernel.
Assume a primitive defined by a splatting kernel $f(x)$ with opacity $o$ and it will be copied $N$ times,  (Bulò et al., 2024; Kheradmand et al., 2024) established that
(12) $o^{\prime}=1-\sqrt[N]{1-o},$
Incorporating $f(x)$ complicates preservation beyond the primitive center, especially at boundaries. Prior approaches like 3DGS-MCMC handled boundary preservation via scale adjustments tailored to Gaussian kernels, which are complex and not directly applicable to deformable kernels.
However, we find that given a regularized small $o$ value, we approximate $o^{\prime}$ using a Taylor expansion $N]{1-o}\approx 1-\frac{o}{N}$, such that
(13) $o^{\prime}=1-\sqrt[N]{1-o}\approx\frac{o}{N}$
Now consider the densified distribution:
(14) $1-\Big{(}1-o^{\prime}\cdot f(x)\Big{)}^{N}\approx 1-\Big{(}1-\frac{o}{N}f(x)\Big{)}^{N}.$
Since $\frac{o}{N}f(x)$ is also small, applying the binomial approximation:
(15) $\Big{(}1-\frac{o}{N}f(x)\Big{)}^{N}\approx 1-N\cdot\frac{o}{N}f(x)+\mathcal{O}(o^{2})=1-o\,f(x)+\mathcal{O}(o^{2}),$
we find:
(16) $1-\big{[}1-o\,f(x)+\mathcal{O}(o^{2})\big{]}=o\,f(x)+\mathcal{O}(o^{2}),$
which matches the original distribution $o\cdot f(x)$.
The error term $\mathcal{O}(o^{2})$ becomes negligible as we employ an opacity regularizer to maintain a small $o$. Therefore, the difference between original distribution and densified distribution diminishes by regularizing opacity regardless of copy times $N$ and splatting kernel $f(x)$. Detailed proof can be found in Figure 7.
Figure 6. Light Decomposition: our Spherical Beta color encoding can effectively decompose diffuse and specular components compared to 3DGS-based approaches using low-order Spherical Harmonics, as illustrated in this figure.
4. Experiments
Methods Mip-NeRF360 Tanks&Temples Deep Blending NeRF Synthetic
[…]
implicit Instant-NGP (Müller et al., 2022) 25.51 0.684 0.398 21.72 0.722 0.330 23.62 0.797 0.423 33.18 0.959 0.055
Mip-NeRF360 (Barron et al., 2021) 27.69 0.792 0.237 22.22 0.758 0.257 29.40 0.900 0.245 33.25 0.962 0.039
Zip-NeRF (Barron et al., 2023) 28.54 0.828 0.189 - - - - - - 33.10 0.971 0.031
[…]
2DGS (Huang et al., 2024) 27.04 0.805 0.297 - - - - - - 33.07 - -
Mip-Splatting (Yu et al., 2023) 27.79 0.827 0.203 - - - - - - 33.33 0.969 0.039
3DGS-MCMC (Kheradmand et al., 2024) 28.29 0.840 0.210 24.29 0.860 0.190 29.67 0.895 0.320 33.80 0.970 0.040
Ours (30k) 28.60 0.844 0.182 24.79 0.868 0.148 30.10 0.910 0.240 34.64 0.973 0.028
Ours (full) 28.75 0.845 0.179 24.85 0.870 0.140 30.12 0.914 0.236 34.66 0.973 0.028
Table 1. Quantitative evaluation was conducted on the Mip-NeRF360 (Barron et al., 2021), Tanks&Temples (Knapitsch et al., 2017), and Deep Blending (Hedman et al., 2018) datasets. We adopt the original number from corresponding papers when available. A dash (“-”) indicates results that were not reported.
We evaluate Deformable Beta Splatting (DBS) on a diverse set of real-world and synthetic datasets to demonstrate its superior visual quality and rendering efficiency compared to existing state-of-the-art methods. Additionally, DBS effectively decouples scene geometry and lighting information, leveraging the learnable frequency characteristics of the Beta Kernel to enhance scene representation.
4.1. Experimental Setup
Implementations.
Our implementation builds upon the 3DGS-MCMC framework (Kheradmand et al., 2024), incorporating modifications to the GSplat library (Ye et al., 2024b). Specifically, we integrate the Beta Kernel and Spherical Beta (SB) functions, which are fully explicit and differentiable, implemented in CUDA for optimized performance. By default, we utilize two light sources in Spherical Beta component per primitive (i.e., $sb=2$), unless otherwise specified. To ensure that the Beta Kernel initially resembles the Gaussian Kernel, we set $b$ to zero at initialization. We observe that different scenes exhibit varying convergence speeds when training the model for the default 30,000 iterations. To accommodate these variations, we have integrated a flexible stopping strategy with patience of 10k iterations into the framework, allowing for two training modes: (Ours, 30K) and (Ours, Full). Detailed training times and iteration counts for each scene are provided in Appendix D. For additional hyperparameters and configuration settings, please refer to Appendix C.
4.2. Results and Comparisons
Radiance Field Reconstruction.
We evaluate DBS using the standard 3DGS evaluation pipeline across several widely recognized datasets, including Mip-NeRF 360 (Barron et al., 2021), Tanks and Temples (Knapitsch et al., 2017), Deep Blending (Hedman et al., 2018), and the NeRF Synthetic dataset (Mildenhall et al., 2021). Performance is measured using standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). For the Mip-NeRF 360, Tanks and Temples, and Deep Blending datasets, we utilize the Colmap Structure-from-Motion (SfM) point cloud provided by the 3DGS (Kerbl et al., 2023) repository for initialization. In contrast, for the NeRF Synthetic dataset, primitives are randomly initialized to simulate diverse starting conditions. We benchmark DBS against prior works, including the state-of-the-art implicit method (ZipNeRF (Barron et al., 2023)) and explicit method (3DGS-MCMC (Kheradmand et al., 2024)). The quantitative results of comparisons are presented in Table 1.
Beta Splatting demonstrates superior performance for radiance field reconstruction across all metrics over almost all datasets compared to prior methods. Both the 30K and Full training modes consistently outperform previous implicit and explicit state-of-the-art methods. See Appendix D for detailed per-scene analysis. Qualitative visual comparisons in Figure 5 reveal that while Zip-NeRF tends to exhibit floater artifacts, and 3DGS-MCMC struggles with rendering sharp specularities, DBS maintains high fidelity in both smooth and highly detailed regions, effectively capturing intricate lighting interactions and geometric details.
Method Mip-NeRF 360
$\text{Training Time}^{\downarrow}$ $\text{Storage Mem}^{\downarrow}$ $\text{Rendering FPS}^{\uparrow}$
Zip-NeRF 5h29m 569.7MB 0.184
3DGS-MCMC 31m 733.19MB 82.46
(Ours,30k,sb=2) 16m 356.04MB 123.11
(Ours,full,sb=2) 22m 356.04MB 123.09
Table 2. Quantitative evaluation of training and rendering efficiency on the Mip-360 dataset
Methods Tanks&Temples Deep Blending
[…]
3DGS, w/ MCMC, sh=3 24.29 437.63 29.67 750.23
DBS, w/ MCMC, sh=3) 24.45 445.05 29.88 762.94
DBS, w/ Kernel-Agnostic MCMC, sh=3 24.57 445.05 30.03 762.94
DBS, w/ Kernel-Agnostic MCMC, sb=1 24.53 155.77 30.04 267.03
DBS, w/ Kernel-Agnostic MCMC, sb=2 24.79 200.27 30.10 343.32
DBS, w/ Kernel-Agnostic MCMC, sb=4 24.81 289.28 30.12 495.91
DBS, w/ Kernel-Agnostic MCMC, sb=6 24.80 378.29 30.08 648.50
[…]
Table 3. Ablation Studies on Tank&Temples and Deep Blending datasets.
Efficiency Analysis.
To evaluate efficiency, we compare our method with Zip-NeRF (Barron et al., 2023) and 3DGS-MCMC (Kheradmand et al., 2024) in terms of training efficiency, rendering efficiency, and model size on a single NVIDIA RTX 6000 Ada Generation GPU.. The summarized efficiency metrics are presented in Table 2, with additional per-scene details available in Appendix D. Our findings indicate that DBS not only achieves better reconstruction quality but also renders 50% faster than 3DGS-MCMC (Kheradmand et al., 2024) under the same number of primitives. When sb=2, our method uses less than half storage memory compared to 3DGS-MCMC while achieving the state-of-the-art in all metrics. These results highlight effectiveness of DBS in delivering high-quality renderings with enhanced computational efficiency.
Ablation Study.
Table 3 evaluates the impact of various configurations and kernel choices on PSNR and memory usage for the Tanks&Temples and Deep Blending datasets. Transitioning from the Gaussian to Deformable Beta kernel in DBS with MCMC improves PSNR from 24.29 to 24.45 and from 29.67 to 29.88, respectively. Adopting Kernel-Agnostic MCMC further increases PSNR to 24.57 and 30.03. Setting the number of spherical beta (sb) light sources to 2 achieves a balanced PSNR of 24.79 and 30.10 with memory usage of 200.27 MB and 343.32 MB. Additionally, incorporating a flexible stopping strategy slightly enhances PSNR to 24.85 while maintaining 30.12 without additional memory cost. Overall, DBS with Kernel-Agnostic MCMC and sb=2 provides the best trade-off between rendering quality and memory consumption, further improved by the flexible stopping strategy.
4.3. Beta Decomposition
Leveraging the deformable nature of the Beta Kernel, our framework enables decomposition of both geometry and lighting post-training. This capability arises from the dual role of the Beta Kernel in encoding volumetric splatting and view-dependent color information.
Geometry Decomposition.
The Beta Kernel’s parameter $b$ intrinsically controls the geometry frequency representation of each primitive. Lower $b$ values correspond to lower-frequency primitives, which predominantly capture the fundamental geometric structures of the scene. Conversely, higher $b$ values encapsulate higher-frequency details, such as textures and fine surface variations.
By setting a mask based on the $b$ parameter, we can selectively isolate primitives based on their frequency contributions. We can also do a simple binary split based on a threshold (e.g. the mean of $b$ values across all primitives). Primitives with $b$ below the threshold primarily represent the scene’s solid geometry, with primitives of $b$ above the threashold provides high-frequency texture details.
Figure 4 illustrates how Beta Kernels separate geometry from texture on real-world datasets. In the ”treehill” scene, the bark texture is represented using primitives with beta parameter $b$. Similarly, in the ”garden” scene, detailed textures on the tabletop are removed, preserving the overall table geometry.
Lighting Decomposition
Since Beta Splatting integrates a learnable Phong Reflection Model (Phong, 1975), we can extracting diffuse component through taking only the base color $c_{0}$. Conversly, only enabling specular component gives the view-dependent specular lighting effects. For comparison, we only enable zero degree of SH for diffuse color while non zero degree for specular effects.
The results in Figure 6 illustrate the effectiveness of this approach. In the ”bicycle” scene, the black metal chair is completely separated from the white highlights, a result unattainable with Gaussian Splatting. Similarly, in the ”bonsai” scene, Beta Splatting successfully reconstructs the bright highlight component of the bonsai container, while Gaussian Splatting fails to capture such extreme view-dependent details.
5. Conclusion
We presented Deformable Beta Splatting (DBS), a novel approach that advances real-time radiance field rendering through three key innovations: Deformable Beta Kernels for adaptive geometry representation, Spherical Betas for efficient view-dependent color encoding, and Kernel-Agnostic MCMC that enhances optimization stability and efficiency by relying solely on regularized opacity. These innovations allow DBS to achieve superior visual quality with less memory and computing resources than prior methods.
Limitations
As our framework is rasterization-based, it occasionally produces popping artifacts due to depth approximation inaccuracies during the sorting process. While adaptive, Spherical Beta Functions struggle to model mirror-like reflections and anisotropic specular highlights effectively. In addition, the frustum model causes Beta Kernels to optimize flat geometries for distant backgrounds, which impacts the overall beta geometry distribution. These limitations offer opportunities for further refinement and enhancement of the framework.
Acknowledgements.
The authors would like to thank our primary sponsors of this research: Mr. Clayton Burford of the Battlespace Content Creation (BCC) team at Simulation and Training Technology Center (STTC). This work is supported by University Affiliated Research Center (UARC) award W911NF-14-D-0005, a Toyota Research Institute University 2.0 program, a gift fund from Dolby, and a gift fund from Google DeepMind. Statements and opinions expressed and content included do not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred.
Supplementary Material
Appendix A 3D Splatting Kernels
A.1. Splatting Condition
In volume rendering and splatting techniques, a continuous 3D kernel function $K(x,y,z)$ is projected onto a 2D plane to produce a continuous 2D kernel function $K(x,y)$. This projection is typically performed along one axis (e.g., the $z$-axis) and is represented by the integral:
[…]
If $K(x,y,z)$ is a separable function, such as:
  • Multiplicative form:
    (18) $K(x,y,z)=K(x)\cdot K(y)\cdot K(z)$
  • Additive form:
    (19) $K(x,y,z)=K(x)+K(y)+K(z)$
It is straightforward to derive the corresponding 2D projection.
However, when we want the kernel to be deformable or to have an adjustable shape (e.g., controlling its spread or sharpness), it may be challenging to compute the integral analytically. For example, consider the radially symmetric kernel:
(20) $K(x,y,z)=1-(x^{2}+y^{2}+z^{2})^{\beta}$
with $\beta>0$ and $x^{2}+y^{2}+z^{2}\leq 1$. Computing the integral to obtain $K(x,y)$ can be difficult due to the non-separable nature of the kernel and the dependence on $\beta$.
A.2. Inverse Splatting
To address this challenge, we can reverse the process by first defining the desired 2D kernel $K(r)$, where $r=\sqrt{x^{2}+y^{2}}$, and then determining whether there exists a corresponding 3D kernel $K(R)$, with $R=\sqrt{x^{2}+y^{2}+z^{2}}$, such that the projection condition is satisfied:
[…]
Assuming radial symmetry and finite support (i.e., $r\leq 1$ and $R\leq 1$), we can utilize the inverse Abel transform to derive the 3D kernel $K(R)$ from the given 2D kernel $K(r)$:
[…]
This approach allows us to define a 2D kernel with the desired properties and ensure that a corresponding 3D kernel exists that satisfies the projection condition.
Example:
Let’s consider the 2D kernel:
(23) $K(r)=(1-r^{2})^{\beta}$
with $\beta>0$ and $0\leq r\leq 1$. The derivative of $K(r)$ is:
(24) $\frac{dK(r)}{dr}=-2\beta r(1-r^{2})^{\beta-1}.$
Substituting into the inverse Abel transform, we obtain the 3D kernel:
[…]
A.3. Requirements for the 2D Kernel Function
To ensure that the 2D kernel $K(r)$ corresponds to a valid 3D kernel $K(R)$ via the inverse Abel transform and satisfies the projection condition, the following requirements must be met:
  1. (1)
    Definition and Range:
    • The kernel is defined for $r\in[0,1]$.
    • The kernel values are in the range $K(r)\in[0,1]$.
  2. (2)
    Boundary Conditions:
    • Opacity at the center:
      (26) $\lim_{r\to 0^{+}}K(r)=1.$
    • Transparency at the boundary:
      (27) $\lim_{r\to 1^{-}}K(r)=0.$
  3. (3)
    Smoothness and Differentiability:
    • The kernel $K(r)$ is continuously differentiable on $(0,1)$:
      (28) $K(r)\in C^{1}(0,1).$
  4. (4)
    Integrability:
    • The kernel and its derivative are absolutely integrable over $(0,1)$:
[…]
  5. (5)
    Existence of Corresponding 3D Kernel:
    • The inverse Abel transform of $K(r)$ yields a valid 3D kernel $K(R)$ that satisfies the projection condition:
[…]
This approach allows us to design kernels with desired properties without needing to compute the 3D kernel explicitly.
Appendix B Kernel-Agnostic Distribution-Preserved Densification
Figure 7. Kernel-Agostic Densification illustration.
Assume a primitive defined by a splatting kernel $f(x)$ with opacity $o$, and it will be cloned $N$ times, Its contribution at a pixel x is given by:
(32) $\alpha=o\cdot f(x)$
When cloning this primitive into $N$ copies, each with new opacity $o^{\prime}$, the combined contribution becomes:
(33) $\alpha^{\prime}=1-(1-o^{\prime}\cdot f(x))^{N}$
To preserve its distribution, we require $\alpha=\alpha^{\prime}$, such that
(34) $o\cdot f(x)=1-(1-o^{\prime}\cdot f(x))^{N}$
Ignoring the kernel momentarily, the condition simplifies to:
(35) $o^{\prime}=1-\sqrt[N]{1-o},$
which aligns with established proof (Bulò et al., 2024; Kheradmand et al., 2024). Incorporating $f(x)$ complicates preservation beyond the primitive center, especially at boundaries. Prior approaches like 3DGS-MCMC handled boundary preservation via scale adjustments tailored to Gaussian kernels, which are complex and not directly applicable to deformable kernels.
We will show that given a small opacity value, only the opacity adjustment is effective and valid regardless the densified number and kernel function.
For small $o$, we approximate $o^{\prime}$ using a Taylor expansion $N]{1-o}\approx 1-\frac{o}{N}$, such that
(36) $o^{\prime}=1-\sqrt[N]{1-o}\approx\frac{o}{N}$
Now consider the densified distribution:
(37) $\alpha^{\prime}=1-\Big{(}1-o^{\prime}\cdot f(x)\Big{)}^{N}\approx 1-\Big{(}1-\frac{o}{N}f(x)\Big{)}^{N}.$
Since $\frac{o}{N}f(x)$ is also small, applying the binomial approximation:
(38) $\Big{(}1-\frac{o}{N}f(x)\Big{)}^{N}\approx 1-N\cdot\frac{o}{N}f(x)+\mathcal{O}(o^{2})=1-o\,f(x)+\mathcal{O}(o^{2}),$
we find:
(39) $\alpha^{\prime}\approx 1-\big{[}1-o\,f(x)+\mathcal{O}(o^{2})\big{]}=o\,f(x)+\mathcal{O}(o^{2}).$
This matches the original distribution:
(40) $\alpha=o\cdot f(x),$
up to first-order in $o$.
The error term $\mathcal{O}(o^{2})$ becomes negligible as $o$ decreases. Therefore, the difference between original distribution and densified distribution diminishes for smaller $o$ regardless of copy times $N$ and splatting kernel $f(x)$.
As illustrated in Figure 7, we can observe that the error decreases as opacity value becomes smaller. When opcacity gets close to 0.1, the two distributions are almost identical.
Appendix C Hyperparameters
The training process is governed by several hyperparameters. The initial learning rate for position optimization is set to $\text{lr}_{\mathbf{\mu}}^{\text{init}}=0.00016$, which decays to a final value of $\text{lr}_{\mathbf{\mu}}^{\text{final}}=0.0000016$. A delay multiplier of $0.01$ is applied, with a maximum number of steps for position learning rate decay of $30{,}000$. The base color learning rate is $\text{lr}_{c_{0}}=0.0025$, and the same value is used for the Spherical Beta parameters learning rate ($\text{lr}_{b_{m},\hat{R}_{m},c_{m}}=0.0025$). The opacity learning rate is relatively high, set to $\text{lr}_{o}=0.05$, while the beta learning rate is $\text{lr}_{b}=0.001$. The scaling and rotation learning rates are $\text{lr}_{\mathbf{s}}=0.005$ and $\text{lr}_{\mathbf{q}}=0.001$, respectively.
Furthermore, the loss function is regularized using a DSSIM loss term with a weight of $\lambda_{\text{SSIM}}=0.2$. Densification occurs every $100$ iterations, starting from $500$ and continuing until $25{,}000$. The background is not randomized during training, and a learning rate noise of $\lambda_{\text{lr}}=5\times 10^{4}$ is used. The regularization terms for scale and opacity are $\lambda_{\Sigma}=0.01$ and $\lambda_{o}=0.01$, respectively.
Appendix D Additional Results
Scenes Zip-NeRF 3DGS-MCMC (Ours,30K,sb=2) (Ours,full,sb=2)
$\text{PSNR}^{\uparrow}$ $\text{SSIM}^{\uparrow}$ $\text{LPIPS}^{\downarrow}$ $\text{PSNR}^{\uparrow}$ $\text{SSIM}^{\uparrow}$ $\text{LPIPS}^{\downarrow}$ $\text{PSNR}^{\uparrow}$ $\text{SSIM}^{\uparrow}$ $\text{LPIPS}^{\downarrow}$ $\text{PSNR}^{\uparrow}$ $\text{SSIM}^{\uparrow}$ $\text{LPIPS}^{\downarrow}$
Mip-NeRF 360 bicycle 25.80 0.769 0.208 26.15 0.81 0.18 26.04 0.804 0.169 26.10 0.805 0.170
flowers 22.40 0.642 0.273 22.12 0.642 0.31 22.44 0.650 0.308 22.49 0.652 0.306
garden 28.20 0.860 0.118 28.16 0.89 0.10 28.15 0.882 0.096 28.11 0.881 0.095
stump 27.55 0.800 0.193 27.80 0.82 0.19 27.56 0.815 0.184 27.50 0.817 0.178
treehill 23.89 0.681 0.242 23.31 0.66 0.29 23.49 0.676 0.293 23.50 0.680 0.289
room 32.65 0.925 0.196 32.48 0.94 0.25 32.83 0.941 0.168 33.03 0.942 0.165
counter 29.38 0.902 0.185 29.51 0.92 0.22 30.36 0.930 0.154 30.70 0.932 0.149
kitchen 32.50 0.928 0.116 32.27 0.94 0.14 32.61 0.939 0.110 32.84 0.940 0.107
bonsai 34.46 0.949 0.173 32.88 0.95 0.22 33.90 0.957 0.156 34.46 0.960 0.151
Mean 28.54 0.828 0.189 28.29 0.84 0.21 28.60 0.844 0.182 28.75 0.845 0.179
[…]
drums 25.84 0.944 0.050 26.29 0.95 0.04 26.78 0.958 0.033 26.81 0.958 0.032
ficus 33.90 0.985 0.015 35.07 0.99 0.01 36.75 0.990 0.010 36.79 0.990 0.010
hotdog 37.14 0.984 0.020 37.82 0.99 0.02 38.85 0.988 0.015 38.95 0.988 0.016
lego 34.84 0.980 0.019 36.01 0.98 0.02 37.12 0.985 0.014 37.11 0.985 0.014
materials 31.66 0.969 0.032 30.59 0.96 0.04 31.12 0.966 0.032 31.13 0.966 0.031
mic 35.15 0.991 0.007 37.29 0.99 0.01 37.66 0.994 0.005 37.67 0.994 0.005
ship 31.38 0.929 0.091 30.82 0.91 0.12 32.13 0.910 0.103 32.15 0.910 0.104
Mean 33.10 0.971 0.031 33.80 0.97 0.04 34.64 0.973 0.028 34.66 0.973 0.028
[…]
train - - - 22.47 0.83 0.24 23.13 0.839 0.185 23.11 0.845 0.174
Mean - - - 24.29 0.86 0.19 24.79 0.868 0.147 24.85 0.873 0.140
[…]
playroom - - - 30.33 0.90 0.31 30.82 0.912 0.240 30.71 0.918 0.236
Mean - - - 29.67 0.90 0.32 30.10 0.910 0.240 30.12 0.914 0.236
Table 4. Quantitative evaluation of rendering results per scene
Metrics bicycle flowers garden stump treehill room counter kitchen bonsai Mean
Number of Primitives 6.0M 3.0M 5.0M 4.5M 3.5M 1.5M 1.5M 1.5M 1.5M 3.1M
[…]
Rendering Time (3DGS-MCMC) 18.77ms 11.53ms 15.38ms 15.75ms 11.33ms 8.27ms 10.37ms 9.09ms 8.62ms 12.12ms
Rendering FPS (3DGS-MCMC) 53.30 86.72 65.00 63.54 88.26 120.86 96.43 110.06 116.01 82.46
Memory (3DGS-MCMC) 1395.42 MB 709.54 MB 1182.56 MB 1064.30 MB 827.79 MB 354.77 MB 354.77 MB 354.77 MB 354.77 MB 733.19 MB
[…]
Training Time (full,sb=2) 24m 11m 25m 14m 12m 19m 32m 28m 32m 22m
Training Steps (full,sb=2) 31.0K 26.5K 25.0K 25.5K 25.5K 48.5K 64.0K 61.5K 84.5 43.5K
Rendering Time (sb=2) 12.70ms 8.13ms 8.71ms 11.33ms 7.42ms 5.03ms 7.69ms 7.04ms 5.04ms 8.12ms
Rendering FPS (sb=2) 78.74 122.93 114.71 88.21 134.82 198.97 130.01 141.93 198.34 123.09
Memory (sb=2) 686.65 MB 343.32 MB 572.21 MB 514.98 MB 400.54 MB 171.66 MB 171.66 MB 171.66 MB 171.66 MB 356.04 MB
Table 5. Quantitative evaluation of training time and rendering efficiency per scene
This detailed per-scene analysis reveals that our method performs comparably to prior works on diffuse scenes, such as ”bicycle,” ”stump,” and ”flowers” from the Mip-NeRF 360 dataset. However, on scenes with complex specular reflections, such as ”counter,” Beta Splatting achieves significant improvements (more than 1 in PSNR) compared to 3DGS-MCMC, Zip-NeRF, and other methods. This highlights the robustness of our approach in capturing intricate light interactions in challenging scenes.